Economics

Suppose that Costa Rican worker Carlos can produce either 100 pounds of coffee or 1 computer per week, and a second worker, Maria, can produce either 150 pounds of coffee or 1 computer per week. Both Carlos and Maria work 50 weeks per year. A third worker, Pedro, joins the Costa Rican economy. Pedro can produce either 140 pounds of coffee or 1 computer per week. Like the other two workers, Pedro works 50 weeks per year.



b. Assume that world prices are such that 1 computer trades for 125 pounds of coffee on international markets and that Costa Rica is open to trade.

 

If coffee can either be produced domestically or imported from abroad in exchange for Costa Rican computers, then the maximum amount of coffee that Carlos, Maria and Pedro can consume is  pounds.  To obtain this amount of coffee, Carlos will produce  , Maria will produce  , and Pedro will produce  .

 

If computers can either be produced domestically or imported from abroad in exchange for Costa Rican coffee, then the maximum number computers that the three workers can consume is  .  To obtain this number of computers, Carlos will produce  , Maria will produce  , and Pedro will produce  .

 

c. Assume instead that world prices are such that 1 computer trades for 200 pounds of coffee on international markets and Costa Rica is open to trade.

 

If coffee can either be produced domestically or imported from abroad in exchange for Costa Rican computers, then the maximum amount of coffee that Carlos and Maria can consume is  pounds.  To obtain this amount of coffee, Carlos will produce  , Maria will produce  , and Pedro will produce  .

 

If computers can either be produced domestically or imported from abroad in exchange for Costa Rican coffee, then the maximum number computers that the three workers can consume is  .  To obtain this number of computers, Carlos will produce  , Maria will produce  , and Pedro will produce  .
Answer #1

Part (c) 2nd question

Since both Carlos and Maria have a comparative advantage in producing coffee with respect to the world prices of computer, they will both produce coffee.

Coffee production (pounds per year)

Carlos: 100*50 = 5000

Maria: 150*50 = 7500

Total coffee production = 12500 pounds

World prices of computers per pound of coffee = 1/80

Maximum number of computers that can be consumed = 12,500 * 1/80 = 156.25

For graph, the 3 points that need to be plotted are:

X coordinate Y coordinate
Computers per year Coffee (pound per year)
100 0
0 12500